A proton, moving with a velocity of viî, collides elastically with another proton that is initially at rest. Assuming that after the collision the speed of the initially moving proton is 2.40 times the speed of the proton initially at rest, find the following.


(a) the speed of each proton after the collision in terms of vi

initially moving proton__ x vi

initially at rest proton__ x vi


(b) the direction of the velocity vectors after the collision (assume that the initially moving proton scatters toward the positive y direction)


initially moving proton ___°relative to the +xdirection

initially at rest proton ___°relative to the +xdirection


1
Expert's answer
2021-11-29T09:15:59-0500
"v_{1f}=\frac{m-m}{m+m}v_{1i}=0\\
v_{2f}=\frac{2m}{m+m}v_{1i}=v_{1i}"